OpenStudy (anonymous):

2n+(7n+8)= missing number n+8

OpenStudy (anonymous):

so is it equal to n+8

OpenStudy (anonymous):

if it is equal to n+8 this is the answer

OpenStudy (anonymous):

Simplifying 2n + (7n + 8) = n + 8 Reorder the terms: 2n + (8 + 7n) = n + 8 Remove parenthesis around (8 + 7n) 2n + 8 + 7n = n + 8 Reorder the terms: 8 + 2n + 7n = n + 8 Combine like terms: 2n + 7n = 9n 8 + 9n = n + 8 Reorder the terms: 8 + 9n = 8 + n Add '-8' to each side of the equation. 8 + -8 + 9n = 8 + -8 + n Combine like terms: 8 + -8 = 0 0 + 9n = 8 + -8 + n 9n = 8 + -8 + n Combine like terms: 8 + -8 = 0 9n = 0 + n 9n = n Solving 9n = n Solving for variable 'n'. Move all terms containing n to the left, all other terms to the right. Add '-1n' to each side of the equation. 9n + -1n = n + -1n Combine like terms: 9n + -1n = 8n 8n = n + -1n Combine like terms: n + -1n = 0 8n = 0 Divide each side by '8'. n = 0 Simplifying n = 0

OpenStudy (anonymous):

no 2n+7n+8 () it the missing number + n+8

OpenStudy (anonymous):

pls show me the real question please

OpenStudy (anonymous):